OpenStudy (anonymous):

Write a system of equations to represent the situation and solve. Adult tickets to the fresno zoo cost 1.50 for students and 2.00 for adults. On one trip, a school bought 56 tickets at a total cost of $107.50. How many people paid adult prices?

OpenStudy (p0sitr0n):

x =number of students 56 -x = adults 1.5 x + 2(56-x) = 107.5 1.5x + 112 - 2x =107.5 -.5x =- 4.5 x =9 56 -x = 47 adults

OpenStudy (anonymous):

Thank you so much!:)